# When Should Selfish Miners Double-Spend?

### v1

@@ 7 Conclusion and Extensions
In this paper, we provided a rigorous and explicit stochastic analysis for the combination of the two celebrated attacks: double-spending and selfish mining. To do so, we first combined stubborn and selfish mining attacks, i.e., constructed a strategy where the attacker acts stubborn until its private branch reaches a certain length $L$, and then switches to act selfish. We provided an analysis of the optimal stubbornness for each parameter regime as well as the maximum stubbornness that is still more profitable than honest mining. Next, we connected the attack to double-spending, as $L>k$ implies double-spending. Thus, given $\alpha$ and $\gamma$, the maximum profitable stubbornness values determine the values of $k$ where $k$-confirmation rule is at risk of double-spending which comes at no-cost to the adversary. We also provided an exact formula for the double-spending probabilities (i.e., the risk) associated with each attack cycle. We provided the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining. In order to conceal the double-spending attack as well as to increase the associated successful double-spending probabilities, we modified the attack that does not match every honest block during the stubborn phase. We provided all the relevant analysis for this modified attack as well.
@@
There are two possible extensions to our paper that can further increase the action space and combining the two would eventually allow all possible state-space considered in [6]. First, note that, we define the attack cycles in a way such that they are i.i.d. This is a design choice which allows us to know whether $tx$ is confirmed or not and use Doob’s optional stopping theorem in our analysis as well as the techniques provided in Appendix A and Appendix B. In this paper, each time an attack cycle starts, we have $A=0$ and $H=0$, since we essentially keep track of the length of the $A$-chain, even if the $H$-chain has some prefix consisting of adversarial blocks of the $A$-chain. On the other hand, one could consider the situations where the duration of attack cycles are defined with respect to the change in the prefix of the $H$-chain, i.e., restricting $H$-chain to consist of the honest blocks exclusively, which is essentially considered in the MDP defined in [6]. This change not only complicates the rigorous analysis and requires us to keep a parameter that tracks the number of honest blocks, but also requires us to keep track of the depth of the $tx$ and $tx^{\prime}$, hence our design choice is to avoid this. The other extension is more obvious, which is to allow trail stubbornness, i.e., to allow the adversary to continue the attack even if it falls behind. By considering an extra parameter for the upper bound on the deficit of the $A$-chain compared to the $H$-chain, our analysis can be redone, however, this would also complicate the analysis provided in Appendix A and Appendix B.
Appendix A Appendix A: Proof of Theorem 3 and Algorithm 1
@@
Let $x_{i}$ and $y_{i}$ be the number of adversarial and honest blocks that make it into the offset chain at the end of an attack cycle $i$. Clearly, the revenue in the long-run is
$\rho=\frac{\sum_{i=1}^{T}x_{i}}{{\sum_{i=1}^{T}x_{i}+\sum_{i=1}^{T}y_{i}}},$ (43)
where $T$ is large. Moving terms, we get $\liminf_{T}\frac{1}{T}\sum_{i}\left((1-\rho)x_{i}-\rho y_{i}\right)=0$, which is the linearization model of MDP in [6]. Next, we show that it can be narrowed down to a specific case of a single cycle, where a decision about overwrite has to be made or delayed, hence we only consider $w_{\rho}(x,y)=(1-\rho)x-\rho y$ and assume $A=H+1=L$.
@@
Let $w^{(L)}_{\rho}(x,y)$ denote the average reward associated with the current cycle of the $L$-stubborn mining strategy when $A=x$ and $H=y$ (cycle has not ended yet necessarily). For example, when $A\geq L$ and $H=A-1$, $w^{(L)}_{\rho}(A,A-1)=(1-\rho)A$ since the decision is to release the $A$-chain immediately. On the other hand, if $A+1=H\leq L$, then $w^{(L)}_{\rho}(H-1,H)=-\rho H$ since the decision is to accept the $H$-chain and abort the attack cycle.
@@
where $w^{(L)}_{\rho_{L}}(\overline{L,L-1})$ is the average reward of $L$-stubborn mining strategy associated with $(\overline{L,L-1})$, i.e., all cycles that do not pass through $A=H+1=L$ Next, for any $\rho\in[0,1]$
@@
Given $A=H+1=L$, the average result associated with $(L+1)$-stubborn mining strategy is related to $L$-stubborn mining strategy in the following way
@@
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage $A-H=2$ implies that the decision will be made when the advantage reduces to $1$, which happens after $\frac{1}{1-2\alpha}$ block arrivals in expectation and $\frac{\alpha}{1-2\alpha}$ of those blocks will be adversarial. Hence, w.p. $\alpha$ we get $w_{\rho}(L+1+\frac{\alpha}{1-2\alpha},0)$.
@@
      The prefix contains $L-i>0$ adversarial blocks w.p. $\gamma(1-\gamma)^{i}$ and we have $w_{\rho}(L-i,i+1)$.
@@
The intermediary results above are now sufficient to prove Theorem 3.
@@
We remark here that for the results of Section 6, we use $10^{-4}$ precision for $\alpha$ and $\gamma$ values and Algorithm 1 runs the while loop at most $4$ times.
Appendix B Appendix B: Proof of Theorem 7 and Algorithm 2
Let $w^{(S)}_{\sigma}(x,y)$ denote the average reward associated with the current cycle of the $S$-stealth mining strategy when $A=x$ and $H=y$.
@@ Appendix B Appendix B: Proof of Theorem 7 and Algorithm 2
00 0 $=\liminf_{T}\frac{1}{T}\sum_{i=1}^{T}w^{(S)}_{\sigma_{S}}(x_{i},y_{i})$ (68)
$=q_{S}w^{(S)}_{\sigma_{S}}(S,S-1)+p_{S}w^{(S)}_{\sigma_{S}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S}),$ (69)
where $w^{(S)}_{\sigma_{S}}(\overline{S})$ is the average reward of $S$-stealth mining strategy associated with $\overline{S}$, i.e., all cycles that do not pass through $A=H+1=S$ or $H=A+1=S$. Next, for any $\sigma\in[0,1]$
$w^{(S)}_{\sigma}(\overline{S})=w^{(S+1)}_{\sigma}(\overline{S}),$ (70)
which is due to the fact that both strategies have the same rewards for all possible arrival paths that do not pass through the event $A=H+1=S$ or $H=A+1=S$ during a cycle. Here, unlike $L$-stubborn mining, we need to exclude $H=A+1=S$ as well. To see this, notice that $(S+1)$-stealth mining does not match the honest block at height $S-1$, whereas $S$-stealth mining does, hence their rewards differ and we cannot apply the separation of dimensions argument explained in Figure 1.
@@
$w^{(S)}_{\sigma_{S}}(\overline{S})<w^{(S+1)}_{\sigma_{S+1}}(\overline{S}).$ (71)
@@
00 0 $=q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S+1)}_{\sigma_{S+1}}(\overline{S})$ (72)
$>q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S})$ (73)
@@
Given $A=H+1=S$, the average result associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
@@
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage $A-H=2$ implies that the decision will be made when the advantage reduces to $1$, hence w.p. $\alpha$ we get $w_{\sigma}(S+1+\frac{\alpha}{1-2\alpha},0)$.
@@
Given $A-1=H=S$, the average result associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
@@
    The prefix contains $S-1$ adversarial blocks w.p. $\gamma$. In this case, we get $w_{\sigma}(S-1,1)$.
@@
by Theorem 7. This in turn, by modifying the arguments of Proposition 1 with the results proven in this section, implies that Algorithm 2 converges to $S^{*}$.

### v2

@@ 7 Conclusion
In this paper, we provided a rigorous and explicit stochastic analysis for the combination of the two celebrated attacks: double-spending and selfish mining. To do so, we first combined stubborn and selfish mining attacks, i.e., constructed a strategy where the attacker acts stubborn until its private branch reaches a certain length $L$, and then switches to act selfish. We provided an analysis of the optimal stubbornness to get the maximum revenue ratio for each parameter regime as well as the maximum stubbornness that is still more profitable than honest mining. Next, we connected the attack to double-spending, as $L>k$ implies double-spending. Thus, given $\alpha$ and $\gamma$, the maximum profitable stubbornness values determine the values of $k$ where $k$-confirmation rule is at risk of double-spending which comes at no-cost to the adversary. We also provided an exact formula for the double-spending probabilities (i.e., the risk) associated with each attack cycle. We provided the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining. In order to conceal the double-spending attack as well as to increase the associated successful double-spending probabilities, we modified the attack that does not match every honest block during the stubborn phase. We provided all the relevant analysis for this modified attack as well.
7.1 Discussion
Notice that, in the literature, the revenue loss of the adversary is usually ignored when considering the double-spending attacks except for the MDP models of [8, 7] where a stochastic explicit analysis is missing and MDPs are used as black boxes to solve the problem. On the other hand, studies such as [16] focus on the double-spend value and only consider the usual double-spending attack structure of [3]. The attack cycles we describe and analyze in this paper allow an attacker to mount a sustained double-spending attack until it succeeds while protecting the attacker from revenue loss in terms of the coinbase rewards in the long run by combining the attack with selfish mining. Here, it is worth noting the severity of the selfish mining and double-spending attacks in general. The selfish mining threatens the fairness of the incentives. A selfish mining attack decreases the revenue and revenue ratio of all honest miners. Under the attack, if the honest mining is not sustainable anymore when the mining costs are taken into account, honest miners leave the system, increasing attackers fraction of hashrate and destabilizing the entire system. Even if the honest mining is sustainable despite the unfair reward distribution, the attacker pool grows by rational miners joining it, destabilizing the entire system again.
On the other hand, the severity of the implications of double-spending attacks depends on how frequent it happens and the value of the transaction at question. It is no doubt that in high-value blockchain systems, double-spending a valuable asset is catastrophic. To prevent such a situation, confirmation latencies are calibrated. For example, with the $k=6$, i.e., the rule-of-the-thumb in Bitcoin, a successful double-spending attack probability for $\alpha=0.1$ is around $10^{-3}$ under $\Delta$-synchrony with $\Delta=10s$, despite the relatively low fork rate compared to for example Litecoin or Ethereum (back when it was PoW) [45]. If we consider a selfish mining attacker with $\alpha=0.3$, the success probability rises above $0.25$. Even if the confirmation latencies are calibrated from $k=6$ to $k=50$ for a high value transaction, the success probability is still around $10^{-4}$. However, these probabilities are obtained assuming worst case scenarios (in the favor of the adversary) in the state-of-the-art models of [45, 44] and do not account for the coinbase reward losses, which potentially disincentivizes the attacker. Our model provides much more practical and realistic model for the attacker since in our model, the attacker is trying to keep an eye on the honest branch and releases its chain if necessary in order not to lose the coinbase rewards.
Even if we limit our focus on selfish mining, our modeling can be useful for analysis of different problems analytically. In our paper, we already solved some, e.g., parameterizing the mining strategies with $L$ and providing a method to find $L^{*}$ that achieves the optimal revenue ratio among the $L$-stubborn mining strategies. Another problem we solved was to parameterize the mining strategies that do not expose the selfish mining attack and to provide a method to find the one that achieves the optimal revenue ratio among these stealth mining strategies. Further, our method is a good analytical substitute for the MDP of [5]. For example, it can be directly used in the analysis of revenue per chain progress for selfish mining attacks considered in [26] where the MDP of [5] cannot be used directly.
From the perspective of consistency of the protocol, $L$-stubborn mining and $S$-stealth mining attacks do not aim to replace the offset chain. The offset chain is extended at the end of every attack cycle and in the long run asymptotically equals the longest chain. At every attack cycle, the aim of the attacker is to cause two parallel forks extending the offset chain, i.e., $A$-chain and $H$-chain. The fork in turn allows the adversary to get a chance at replacing the honest blocks of $H$-chain, stealing their revenue and double-spending. The fork that loses this competition at the end of each attack cyle essentially joins the orphan blocks of the blockchain tree. Within this attack model, the block at the tip of the offset chain can be seen as a Nakamoto block [43], i.e., it is permanently confirmed since both the attacker and the honest miner accept them. On the other hand, double-spending happens because the honest miners and merchants accept the blocks according to the $k$-confirmation rule which is a heuristic and are not able to pinpoint the tip of the offset chain, especially if the attacker is using $S$-stealth mining attack to conceal its $A$-chain by avoiding the matching action.
7.2 Extensions
@@
There are two possible extensions to our paper that can further increase the action space and combining the two would eventually allow all possible state-space considered in [5]. First, note that, we define the attack cycles in a way such that they are i.i.d. This is a design choice which allows us to know whether $tx$ is confirmed or not and use Doob’s optional stopping theorem in our analysis as well as the techniques provided in Appendix B and Appendix C. In this paper, each time an attack cycle starts, we have $A=0$ and $H=0$, since we essentially keep track of the length of the $A$-chain, even if the $H$-chain has some prefix consisting of adversarial blocks of the $A$-chain. On the other hand, one could consider the situations where the duration of attack cycles are defined with respect to the change in the prefix of the $H$-chain, i.e., restricting $H$-chain to consist of the honest blocks exclusively, which is essentially considered in the MDP defined in [5]. This change not only complicates the rigorous analysis and requires us to keep a parameter that tracks the number of honest blocks, but also requires us to keep track of the depth of the $tx$ and $tx^{\prime}$, hence our design choice is to avoid this. The other extension is more obvious, which is to allow trail stubbornness, i.e., to allow the adversary to continue the attack even if it falls behind. By considering an extra parameter for the upper bound on the deficit of the $A$-chain compared to the $H$-chain, our analysis can be redone, however, this would also complicate the analysis provided in Appendix B and Appendix C. Note that, increasing the action space for the attacker increases the maximum revenue ratio that can be obtained. However, based on our simulations and comparison with the maximum revenue ratios obtained by the MDP of [5] for several $(\alpha,\gamma)$ parameters, the difference is negligible.
Appendix A Selfish Mining [4] and OSM MDP [5]
Here, we restate and summarize the important bits of the selfish mining attack of [4] as well as the Optimal Selfish Mining (OSM) MDP model introduced in [5] to find the $\epsilon$-optimal selfish mining strategies.
A.1 Selfish Mining of Eyal and Sirer [4]
In essence, the original attack introduced in [4] works as follows. Stating from the same offset chain:
  1. 1.
    If the honest miners mine a block first, the attacker accepts the block and redefines the offset chain.
  2. 2.
    If the attacker pool mines a block first, it hides it and tries to extend it. When the next block arrives:
    1. (a)
      If the block is adversarial, the attacker has $2$ private blocks. It will keep mining privately until the difference between the public branch and private branch reduces to $1$, at which point it will release the private branch to override the public branch.
    2. (b)
      If the block is honest, the adversary releases its private branch, i.e., matching, resulting in a public fork of length $1$. At this point, one branch is adversarial and the other is honest, and $\gamma$ fraction of the honest miners prefer the adversarial branch to mine on. Whoever mines the next block, releases it immediately, and the offset chain is redefined.
Notice that the attacker avoids the release of its block at Step 2 initially. Our idea is to extend this further even in Step 2a if the adversarial branch is not long enough, i.e., until the private branch reaches length $L$. This is because, an override action means that the attacker is releasing one additional block to convince others that its released branch is longer than the honest branch, however, it also means the adversary is potentially losing some additional revenue.
A.2 OSM MDP of [5]
We start by stating the state space of the OSM MDP. A state is essentially of the $3$-tuple form $(a,h,fork)$, where $a$ and $h$ denote the lengths of the adversarial private and honest public branches, respectively. The last field forkfork stores the information about the feasibility of the match action. If $fork=relevant$, it means the adversary can take the action whereas if $fork\neq relevant$, it cannot. On the other hand, if a match action is taken recently, $fork=active$. We summarize the state action space and the transitions in Table 2, taken directly from [5].
Table 2: State, action, and transition of OSM MDP [5]
State $\times$ Action State Probability Reward
$(a,h,\cdot),\text{adopt}$ $(1,0,\text{irrelevant})$ $\alpha$ $(0,h)$
$(0,1,\text{irrelevant})$ $1-\alpha$ $(0,h)$
$(a,h,\cdot),\text{override}^{\dagger}$ $(a-h,0,\text{irrelevant})$ $\alpha$ $(h+1,0)$
$(a-h-1,1,\text{relevant})$ $1-\alpha$ $(h+1,0)$
$(a,h,\text{irrelevant}),\text{wait}$ $(a+1,h,\text{irrelevant})$ $\alpha$ $(0,0)$
$(a,h,\text{relevant}),\text{wait}$ $(a,h+1,\text{relevant})$ $1-\alpha$ $(0,0)$
$(a,h,\text{active}),\text{wait}$ $(a+1,h,\text{active})$ $\alpha$ $(0,0)$
$(a,h,\text{relevant}),\text{match}^{\ddagger}$ $(a-h,1,\text{relevant})$ $\gamma\cdot(1-\alpha)$ $(h,0)$
$(a,h+1,\text{relevant})$ $(1-\gamma)\cdot(1-\alpha)$ $(0,0)$
$\ddagger$ feasible only when $a\geq h$
$\dagger$ feasible only when $a>h$
Letting $r^{t}=(r^{t}_{1},r^{t}_{2})$ denote the reward for the transitions between the states according to the action taken at step $t$, we get the adversarial revenue ratio as
$\rho$ $=\mathbb{E}\left[\liminf_{T\rightarrow\infty}\frac{\sum_{t=1}^{T}r^{t}_{1}}{{\sum_{t=1}^{T}r^{t}_{1}+\sum_{t=1}^{T}r^{t}_{2}}}\right].$ (43)
Next, define $w_{\rho}(x,y)$ as
$w_{\rho}(x,y)=(1-\rho)x-\rho y.$ (44)
Now, for any policy $\pi$, if (43) is rearranged, we get
$v^{\pi}_{\rho}$ $=\mathbb{E}\left[\liminf_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}w_{\rho}(r^{t}_{1},r^{t}_{2})\right].$ (45)
If a policy $\pi^{*}$ maximizes the revenue ratio and results in $\rho^{*}$, then $v^{\pi^{*}}_{\rho^{*}}=0$. Hence, [5] uses the monotonicity of $w_{\rho}(x,y)$ in $\rho$ to find such an optimal policy. For further details about the OSM MDP formulation and its analysis, we refer the reader to [5].
Appendix B Proof of Theorem 2 and Algorithm 1
As in [5], we define $w_{\rho}(x,y)$ as
$w_{\rho}(x,y)=(1-\rho)x-\rho y.$ (46)
@@ Appendix B Proof of Theorem 2 and Algorithm 1
Let $x_{i}$ and $y_{i}$ be the number of adversarial and honest blocks that make it into the offset chain at the end of an attack cycle $i$ of $L$-stubborn mining strategy. Clearly, the revenue ratio in the long run is
$\rho$ $=\frac{\sum_{i=1}^{T}x_{i}}{{\sum_{i=1}^{T}x_{i}+\sum_{i=1}^{T}y_{i}}}$ (47)
$=\frac{\mathbb{E}[x_{i}]}{\mathbb{E}[x_{i}+y_{i}]},$ (48)
where $T$ is large and (48) follows from the stopping time, i.e., the attack cycles are i.i.d. Moving terms, we get $w_{\rho}(\mathbb{E}[x_{i}],\mathbb{E}[y_{i}])=0$, which is similar to the linearization of [5].
Notice that, on average, $\rho$ fraction of the rewards belongs to the adversary. Hence, $w_{\rho}(x_{i},y_{i})$ is essentially a function that indicates how much the reward distribution of the cycle $i$ deviates from the average. Using this interpretation, we show that, when the strategies are restricted to $L$-stubborn mining strategies, the analysis can be narrowed down to a specific case of a single cycle $(A=H+1)$, where a decision about override has to be made or delayed. Here, our goal is to prove that, we can only check the cases of $A=H+1$, instead of implementing an MDP directly.
Assume, the current cycle $i$ of the $L$-stubborn mining strategy has $A=x$ and $H=y$ (cycle has not ended yet necessarily). Let $w^{(L)}_{\rho}(x,y)$ denote the expected value of $w_{\rho}$ associated with the cycle $i$ given $A=x$ and $H=y$ and the attacker employs $L$-stubborn mining strategy, i.e.,
$w^{(L)}_{\rho}(x,y)$ $=w_{\rho}(\mathbb{E}^{(L)}[x_{i}|A=x,H=y],\mathbb{E}^{(L)}[y_{i}|A=x,H=y])$ (49)
$=(1-\rho)\mathbb{E}^{(L)}[x_{i}|A=x,H=y]-\rho\mathbb{E}^{(L)}[y_{i}|A=x,H=y],$ (50)
@@
where the superscript $(L)$ means that the adversary follows $L$-stubborn mining strategy. For example, when $A\geq L$ and H=A1H=A-1, $w^{(L)}_{\rho}(A,A-1)=(1-\rho)A$ since the decision is to release the $A$-chain immediately. On the other hand, if $A+1=H\leq L$, where $H$-chain has $a$ adversarial blocks in its prefix, then $w^{(L)}_{\rho}(H-1,H)=(1-\rho)a-\rho(H-a)$ since the decision is to accept the $H$-chain and abort the attack cycle.
@@
where $w^{(L)}_{\rho_{L}}(\overline{L,L-1})$ is the expected value of $w_{\rho}$ for $L$-stubborn mining strategy given $(\overline{L,L-1})$, i.e., the cycle does not pass through $A=H+1=L$. Next, for any $\rho\in[0,1]$
@@
Given $A=H+1=L$, the expected result of $w_{\sigma}$ associated with $(L+1)$-stubborn mining strategy is related to $L$-stubborn mining strategy in the following way
@@
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage AH=2A-H=2 implies that the decision will be made when the advantage reduces to $1$, which happens after $\frac{1}{1-2\alpha}$ block arrivals in expectation and $\frac{\alpha}{1-2\alpha}$ of those blocks will be adversarial. Hence, w.p. $\alpha$ we get $w_{\rho}(L+1+\frac{\alpha}{1-2\alpha},0)$.
@@
      The prefix contains Li>0L-i>0 adversarial blocks w.p. $\gamma(1-\gamma)^{i}$ and we have $w_{\rho}(L-i,i+1)$.
@@
Notice what we did in this sequence of lemmas: In Lemma 3, given the current attack cycle has $A=L=H+1$, we established a relation between the expected value of $w_{\rho_{L}}$ and $w_{\rho_{L+1}}$ for $L$-stubborn mining strategy and $L+1$-stubborn mining strategy assuming $\rho_{L}>\rho_{L+1}$ as well as the opposite situation in Corollary 1. On the other hand, in Lemma 4, we established a direct relation between the expected value of $w_{\rho}$ for $L$-stubborn mining strategy and $L+1$-stubborn mining strategy. Hence, to prove Theorem 2, all that is left is to check when these relationships hold.
@@
We remark here that for the results of Section 6, we use $10^{-4}$ precision for $\alpha$ and $\gamma$ values and Algorithm 1 runs the while loop at most 44 times.
Appendix C Proof of Theorem 6 and Algorithm 2
Let $w^{(S)}_{\sigma}(x,y)$ denote the expected value of $w_{\sigma}$ associated with the cycle $i$ given $A=x$ and $H=y$ and the attacker employs $S$-stealth mining strategy, i.e.,
$w^{(S)}_{\sigma}(x,y)$ $=w_{\sigma}(\mathbb{E}^{(S)}[x_{i}|A=x,H=y],\mathbb{E}^{(S)}[y_{i}|A=x,H=y])$ (74)
$=(1-\sigma)\mathbb{E}^{(S)}[x_{i}|A=x,H=y]-\sigma\mathbb{E}^{(S)}[y_{i}|A=x,H=y].$ (75)
@@ Appendix C Proof of Theorem 6 and Algorithm 2
$0$ $=\liminf_{T}\frac{1}{T}\sum_{i=1}^{T}w^{(S)}_{\sigma_{S}}(x_{i},y_{i})$ (77)
$=q_{S}w^{(S)}_{\sigma_{S}}(S,S-1)+p_{S}w^{(S)}_{\sigma_{S}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S-1}),$ (78)
where $w^{(S)}_{\sigma_{S}}(\overline{S-1})$ is the the expected value of $w_{\rho}$ for $S$-stealth mining strategy given $\overline{S-1}$, i.e., the cycle does not pass through $A=H+1=S$ or $H=A+1=S$. Next, for any $\sigma\in[0,1]$
$w^{(S)}_{\sigma}(\overline{S-1})=w^{(S+1)}_{\sigma}(\overline{S-1}),$ (79)
which is due to the fact that both strategies have the same rewards for all possible arrival paths that do not pass through the event $A=H+1=S$ or $H=A+1=S$ during a cycle. Here, unlike $L$-stubborn mining, we need to exclude $H=A+1=S$ as well. To see this, notice that $(S+1)$-stealth mining does not match the honest block at height S1S-1, whereas $S$-stealth mining does, hence their rewards differ and we cannot apply the separation of dimensions argument explained in Figure 1.
@@
$w^{(S)}_{\sigma_{S}}(\overline{S-1})<w^{(S+1)}_{\sigma_{S+1}}(\overline{S-1}).$ (80)
@@
$0$ $=q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S+1)}_{\sigma_{S+1}}(\overline{S-1})$ (81)
$>q_{S}w^{(S+1)}_{\sigma_{S+1}}(S,S-1)+p_{S}w^{(S+1)}_{\sigma_{S+1}}(S-1,S)+(1-q_{S}-p_{S})w^{(S)}_{\sigma_{S}}(\overline{S-1})$ (82)
@@
Given $A=H+1=S$, the expected value of $w_{\sigma}$ associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
@@
    The next block is adversarial w.p. $\alpha$. In this case, the adversarial advantage AH=2A-H=2 implies that the decision will be made when the advantage reduces to $1$, hence w.p. $\alpha$ we get $w_{\sigma}(S+1+\frac{\alpha}{1-2\alpha},0)$.
@@
Given A1=H=SA-1=H=S, the expected value of $w_{\sigma}$ associated with $(S+1)$-stealth mining strategy is related to $S$-stealth mining strategy in the following way,
@@
    The prefix contains S1S-1 adversarial blocks w.p. $\gamma$. In this case, we get $w_{\sigma}(S-1,1)$.
@@
by Theorem 6. This in turn, by modifying the arguments of Proposition 1 with the results proven in this section, implies that Algorithm 2 returns $S^{*}$.